### v1

# Eugene Kritchevski

Abstract
We introduce an elementary framework that captures the mechanism driving the alignment of rows and columns in products of positive matrices. All worst-case misalignment occurs already in dimension two, leading to an explicit collapse principle and a sharp nonlinear bound for finite products. The proof avoids Hilbert-metric and cone-theoretic techniques, relying instead on basic calculus. In the Hilbert metric, the classical Birkhoff–Bushell contraction captures only the linearized asymptotic regime, whereas our nonlinear envelope function gives the exact worst-case behavior for finite products.
1 Introduction
Products of positive matrices exhibit a pronounced alignment phenomenon: as the product grows, all rows and columns become nearly proportional. Classical methods approach this behavior through the Hilbert projective metric, probabilistic tools, or spectral theory. In this work we take a different route: we work directly with the distortion
$R(A):=\max_{i,j,k,\ell}\frac{a_{ik}a_{j\ell}}{a_{i\ell}a_{jk}},\qquad A\in\mathbb{R}_{>0}^{d_{1}\times d_{2}},$
which quantifies how far the rows or columns of $A$ are from being proportional. We have $R(A)=1$ in the perfectly aligned case; in general $R(A)\geq 1$ and increases with misalignment. Although $R(A)$ corresponds to the exponential of the projective diameter of $A$, we rely only on its elementary ratio form and use no projective-metric theory.
Tracking how distortion changes under matrix multiplication gives an elementary, coordinate-level mechanism for alignment. The central observation is that worst–case distortion occurs already in dimension two; higher dimensions do not produce any new extremal behaviour. In dimension two, the explicit sharp distortion bound can then be computed by an elementary one–variable calculus optimization. A dimension collapse argument shows that exactly the same bound holds in arbitrary dimension. This yields the envelope inequality
$R(AB)\leq\Phi(R(A),R(B)),$
where
$\Phi(\alpha,\beta)=\left(\frac{1+\sqrt{\alpha\beta}}{\sqrt{\alpha}+\sqrt{\beta}}\right)^{\!2}.$
Roadmap.
Section 2 introduces distortion and proves the envelope inequality in dimension 2. Section 3 extends the envelope inequality to arbitrary dimension via a collapse argument. Section 4 develops distortion for longer products of matrices, including explicit Frobenius–Perron bounds. Section 5 compares our result with the classical Birkhoff–Bushell theory. Section 6 highlights several directions for further work. Additional calculus lemmas used in the proofs are collected in the Appendix.
2 Distortion Calculus in Dimension 22
Definition 2.1.
Let
$A=\begin{pmatrix}a_{11}&a_{12}\\[2.0pt] a_{21}&a_{22}\end{pmatrix}$
be a $2\times 2$ matrix with strictly positive entries.
  1. 1.
    The oriented distortion of $A$ is
    $F(A):=\frac{a_{11}a_{22}}{\,a_{12}a_{21}\,}.$
  2. 2.
    The distortion of $A$ is
    $R(A):=\max\{F(A),\,1/F(A)\}.$
Both $F(A)$ and $R(A)$ measure how far the two columns (equivalently, the two rows) of $A$ are from being proportional. Perfect proportionality is equivalent to $F(A)=R(A)=1$. The following properties are immediate from the definition:
  • Scaling invariance. Multiplying any row or any column of $A$ by a positive scalar leaves both $F(A)$ and $R(A)$ unchanged.
  • Effect of swapping. Swapping the two rows or the two columns of $A$ replaces $F(A)$ by $1/F(A)$ and therefore leaves $R(A)$ unchanged.
  • Orientation via determinant. For $2\times 2$ positive matrices,
    $R(A)=F(A)\iff F(A)\geq 1\iff\det(A)\geq 0,$
    and
    $R(A)=1/F(A)\iff F(A)\leq 1\iff\det(A)\leq 0.$
    In the perfectly aligned case, $\det(A)=0$ and both orientations coincide. Note that in our projective, scale-invariant setting, the area interpretation of the determinant is irrelevant; only the sign information matters.
Our goal is to obtain a sharp explicit upper bound on the distortion
$R(AB)$
of the product of two positive $2\times 2$ matrices. The analysis is greatly simplified by a few distortion-preserving normalizations.
We first ensure positive orientation:
$F(A)\geq 1,\qquad F(B)\geq 1,$
so that $F(A)=R(A)$ and $F(B)=R(B)$. This can always be achieved by swapping the rows of $A$ and/or the columns of $B$; such swaps do not affect the distortion of $A$, $B$, or $AB$. After ensuring positive orientation of $A$ and $B$, we have $\det(A)\geq 0$, $\det(B)\geq 0$, and so $\det(AB)=\det(A)\det(B)\geq 0$. Then $F(AB)\geq 1$ and $AB$ is also positively oriented: $F(AB)=R(AB)$.
Next, we are free to multiply the rows of $A$ and the columns of $B$ by any positive scalars. Such scalings affect neither the orientation nor the distortion of $A$, $B$, or $AB$.
After these normalizations, we may assume, without loss of generality, that $A$ and $B$ are in the canonical form
$A=\begin{pmatrix}1&u\\[5.69054pt] 1&\alpha u\end{pmatrix},\qquad B=\begin{pmatrix}1&1\\[5.69054pt] v&\beta v\end{pmatrix}$,
where
$\alpha=R(A)\geq 1,\qquad\beta=R(B)\geq 1.$
A direct computation gives
$AB=\begin{pmatrix}1+uv&1+\beta\,uv\\[5.69054pt] 1+\alpha\,uv&1+\alpha\beta\,uv\end{pmatrix}=\begin{pmatrix}1+t&1+\beta t\\[5.69054pt] 1+\alpha t&1+\alpha\beta t\end{pmatrix}$,
where
$t=uv>0.$
Then
$R(AB)=f_{\alpha,\beta}(uv),$
where
$f_{\alpha,\beta}(t):=\frac{(1+t)(1+\alpha\beta t)}{(1+\beta t)(1+\alpha t)}\qquad t>0.$
Thus the problem of finding a sharp upper bound for $R(AB)$ in terms of $R(A)$ and $R(B)$ reduces to a simple calculus exercise: it suffices to maximize the function $f_{\alpha,\beta}(t)$ for $t>0$. The details are carried out in Lemma A.1 and the result is
$\max_{t>0}f_{\alpha,\beta}(t)=\Phi(\alpha,\beta),$
where
$\Phi(\alpha,\beta)=\frac{\bigl(1+\sqrt{\alpha\beta}\bigr)^{2}}{\bigl(\sqrt{\alpha}+\sqrt{\beta}\bigr)^{2}}.$
We refer to $\Phi$ as the envelope function.
Thus we obtain
$R(AB)=f_{\alpha,\beta}(uv)\leq\Phi(\alpha,\beta),$
which proves the following bound:
Theorem 2.2 (Envelope Inequality in Dimension $2$).
Let $A$ and $B$ be strictly positive $2\times 2$ matrices. Then
$\boxed{R(AB)\leq\Phi\bigl(R(A),R(B)\bigr)=\left(\frac{1+\sqrt{R(A)R(B)}}{\sqrt{R(A)}+\sqrt{R(B)}}\right)^{\!2}}.$
Moreover, the envelope function $\Phi(\alpha,\beta)$ is increasing in each variable (Lemma A.2). Therefore, from Theorem 2.2 we obtain:
Corollary 2.3.
Let $A$ and $B$ be strictly positive $2\times 2$ matrices, and suppose
$R(A)\leq\alpha,\qquad R(B)\leq\beta.$
Then
$R(AB)\leq\Phi(\alpha,\beta).$
We note that the envelope inequality is not merely an upper bound; the maximal distortion is attainable.
Proposition 2.4 (Two–Sided Sharpness).
Let $\alpha,\beta\geq 1$.
  1. (i)
    Given any $A>0$ with $R(A)=\alpha$, one can choose $B>0$ with $R(B)=\beta$ such that
    $R(AB)=\Phi(\alpha,\beta).$
  2. (ii)
    Given any $B>0$ with $R(B)=\beta$, one can choose $A>0$ with $R(A)=\alpha$ such that
    $R(AB)=\Phi(\alpha,\beta).$
Proof.
We first prove (i). By Lemma A.1, the function $f_{\alpha,\beta}(t)$ attains its maximum $\Phi(\alpha,\beta)$ at
$t^{*}=\frac{1}{\sqrt{\alpha\beta}}.$
Every product of the form
$\begin{pmatrix}1&u\\[5.69054pt] 1&\alpha u\end{pmatrix}\begin{pmatrix}1&1\\[5.69054pt] v&\beta v\end{pmatrix},\qquad uv=t^{*},$
has distortion $\Phi(\alpha,\beta)$.
Let $A>0$ be given with $R(A)=\alpha$. By rescaling the rows of $A$ by positive factors (which does not change $R(A)$ nor $R(AB)$), we may assume that
$A=\begin{pmatrix}1&x\\[5.69054pt] 1&y\end{pmatrix},\qquad x,y>0.$
If $A$ is positively oriented, i.e. $y/x=\alpha$, we take
$B=\begin{pmatrix}1&1\\[5.69054pt] v&\beta v\end{pmatrix},\qquad v=\frac{t^{*}}{x}.$
If $A$ is negatively oriented, i.e. $x/y=\alpha$, we take
$B=\begin{pmatrix}1&1\\[5.69054pt] \beta v&v\end{pmatrix},\qquad v=\frac{t^{*}}{y}.$
In both cases one checks directly that
$R(AB)=\Phi(\alpha,\beta).$
This proves (i).
For (ii), use that $R(M)=R(M^{\mathsf{T}})$ and $\Phi(\alpha,\beta)=\Phi(\beta,\alpha)$. Given $B$ with $R(B)=\beta$, apply (i) to $B^{\mathsf{T}}$ and $\alpha$: there exists $C>0$ with $R(C)=\alpha$ such that
$R(B^{\mathsf{T}}C)=\Phi(\beta,\alpha).$
Setting $A=C^{\mathsf{T}}$ gives
$R(AB)=R((AB)^{\mathsf{T}})=R(B^{\mathsf{T}}C)=\Phi(\beta,\alpha)=\Phi(\alpha,\beta),$
which proves (ii). ∎
Corollary 2.5 (optimization identity).
For all $\alpha,\beta\geq 1$,
$\boxed{\Phi(\alpha,\beta)=\max\bigl\{\,R(AB):A,B>0,\ R(A)\leq\alpha,\ R(B)\leq\beta\,\bigr\}}.$
3 Reduction to Two Dimensions in the General Case
The goal of this section is to extend the envelope inequality to arbitrary dimension.
For $x,y\in\mathbb{R}_{>0}^{d}$, define
$s^{+}(x,y)=\max_{i}\frac{y_{i}}{x_{i}},\qquad s^{-}(x,y)=\min_{i}\frac{y_{i}}{x_{i}},$
and the distortion
$\operatorname{Dist}(x,y)=\frac{s^{+}(x,y)}{s^{-}(x,y)}.$
For perfectly aligned vectors we have $s^{+}(x,y)=s^{-}(x,y)$, so $\operatorname{Dist}(x,y)=1$. When the ratios $y_{i}/x_{i}$ are more scattered, the distortion increases. Figure 1 illustrates $s^{+}(x,y)$ and $s^{-}(x,y)$ as maximal and minimal slopes of the points $(x_{i},y_{i})$ in $\mathbb{R}_{>0}^{2}$.
Figure 1: The pair $(x,y)$ represented as a cloud of points between the extremal rays of slopes $s^{+}(x,y)$ and $s^{-}(x,y)$ in $\mathbb{R}_{>0}^{2}$.
We extend the definition of distortion to positive matrices of size $d_{1}\times d_{2}$ by
$R(A)=\max_{\begin{subarray}{c}1\leq i,j\leq d_{1}\\ 1\leq k,\ell\leq d_{2}\end{subarray}}\frac{a_{ik}\,a_{j\ell}}{a_{i\ell}\,a_{jk}}.$
Thus $R(A)$ is the maximum of the distortions of all $2\times 2$ submatrices of $A$, obtained by choosing any two rows and any two columns. Equivalently, it is the maximum distortion between any two columns (or two rows) of $A$:
$R(A)=\max_{k,\ell}\operatorname{Dist}\bigl(A_{*k},A_{*\ell}\bigr)=\max_{i,j}\operatorname{Dist}\bigl(A_{i*},A_{j*}\bigr).$
Define $F_{d}:(\mathbb{R}^{d}_{>0})^{4}\to\mathbb{R}_{>0}$ by
$F_{d}(x,y;u,v):=\frac{(x\cdot u)(y\cdot v)}{(x\cdot v)(y\cdot u)}.$
Here $x\cdot u$ denotes the standard Euclidean dot product.
Lemma 3.1 (coordinatewise extremal ratio pushing).
Let $d\geq 2$ and let $x,y,u,v\in\mathbb{R}^{d}_{>0}$. Define
$a^{\pm}:=s^{\pm}(x,y),\qquad b^{\pm}:=s^{\pm}(u,v).$
Then there exist vectors $y^{*},v^{*}\in\mathbb{R}^{d}_{>0}$ such that
$F_{d}(x,y^{*};u,v^{*})\;\geq\;F_{d}(x,y;u,v),$
$s^{\pm}(x,y^{*})=s^{\pm}(x,y),\qquad s^{\pm}(u,v^{*})=s^{\pm}(u,v),$
and for all $i$,
$\frac{y_{i}^{*}}{x_{i}}\in\{a^{-},a^{+}\},\qquad\frac{v_{i}^{*}}{u_{i}}\in\{b^{-},b^{+}\}.$
Proof.
We will first construct $y^{*}$ and then use a similar procedure to obtain $v^{*}$. The strategy is to adjust the coordinates of $y$ one at a time, always modifying them in a way that does not decrease the value of $F_{d}(x,y;u,v)$, and while keeping the extremal slopes $s^{\pm}(x,y)$ unchanged.
To make this precise, partition the index set $\{1,\dots,d\}$ into
$I_{-}:=\{\,i:y_{i}/x_{i}=a^{-}\,\},\qquad I_{+}:=\{\,i:y_{i}/x_{i}=a^{+}\,\},$
and
$I_{0}:=\{1,\dots,d\}\setminus(I_{-}\cup I_{+}).$
We will never alter the coordinates in $I_{-}$ or $I_{+}$, so that the extremal slopes remain fixed at all times. Thus we immediately set
$y_{i}^{*}:=y_{i}\qquad\text{for all }i\in I_{-}\cup I_{+}.$
The coordinatewise replacement procedure is applied only to indices in $I_{0}$.
We begin with a simple coordinate monotonicity observation. Fix $x,u,v$ and all coordinates of $y$ except $y_{i}$. Then
$f(y_{i}):=F_{d}(x,y;u,v)$
is a fractional linear function of $y_{i}$ with strictly positive denominator, and is therefore monotone (either increasing or decreasing) in $y_{i}$. If $f$ is increasing in $y_{i}$, we set
$y_{i}^{*}:=a^{+}\,x_{i},$
while if $f$ is decreasing in $y_{i}$, we set
$y_{i}^{*}:=a^{-}\,x_{i}.$
Then
$F_{d}(x,y^{*};u,v)\geq F_{d}(x,y;u,v),\qquad s^{\pm}(x,y^{*})=s^{\pm}(x,y),$
and
$\frac{y_{i}^{*}}{x_{i}}\in\{\,a^{-},\,a^{+}\,\}.$
The key point is that replacing $(x_{i},y_{i})$ by $(x_{i},y_{i}^{*})$ does not decrease $F_{d}$, does not change $s^{\pm}(x,y)$, and forces the ratio $y_{i}^{*}/x_{i}$ to be extremal. The frozen extremal indices in $I_{-}\cup I_{+}$ ensure that the original extremal ratios are preserved throughout.
We now apply this replacement successively to all $i\in I_{0}$, always leaving previously modified coordinates unchanged. This yields a vector $y^{*}$ such that
$s^{\pm}(x,y^{*})=s^{\pm}(x,y),\qquad F_{d}(x,y^{*};u,v)\geq F_{d}(x,y;u,v),$
and every ratio $y_{i}^{*}/x_{i}$ is extremal:
$\frac{y_{i}^{*}}{x_{i}}\in\{\,a^{-},\,a^{+}\,\}.$
With $y^{*}$ fixed, we now apply exactly the same coordinatewise extremal-pushing procedure to the pair $(u,v)$. We freeze the indices where the ratios $v_{i}/u_{i}$ already attain the extremal values $b^{\pm}$, and modify only the remaining coordinates exactly as above: for each active coordinate we replace $v_{i}$ by either $b^{-}u_{i}$ or $b^{+}u_{i}$, according to the monotonicity of $F_{d}(x,y^{*};u,v)$ in $v_{i}$. This produces a vector $v^{*}$ such that
$s^{\pm}(u,v^{*})=s^{\pm}(u,v),\qquad F_{d}(x,y^{*};u,v^{*})\geq F_{d}(x,y^{*};u,v),$
and each ratio $v_{i}^{*}/u_{i}$ is extremal:
$\frac{v_{i}^{*}}{u_{i}}\in\{\,b^{-},\,b^{+}\,\}.$
This completes the proof. ∎
Lemma 3.2 ($2$-point support collapse).
Assume that the coordinate ratios of $y^{*}/x$ take only the two extremal values $a^{\pm}$, and the coordinate ratios of $v^{*}/u$ take only the two extremal values $b^{\pm}$, where
$a^{\pm}=s^{\pm}(x,y^{*}),\qquad b^{\pm}=s^{\pm}(u,v^{*}).$
Then there exist vectors $x^{\prime},y^{\prime},u^{\prime},v^{\prime}\in\mathbb{R}^{2}_{>0}$ such that
$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})\;=\;F_{d}(x,y^{*};u,v^{*}),$
and
$s^{\pm}(x^{\prime},y^{\prime})=a^{\pm},\qquad s^{\pm}(u^{\prime},v^{\prime})=b^{\pm}.$
Proof.
If $a^{+}=a^{-}=a$, then $y^{*}=ax$ and hence $F_{d}(x,y^{*};u,v^{*})=1$. The lemma then holds, for example, with
$x^{\prime}=(1,1),\qquad y^{\prime}=(a,a),\qquad u^{\prime}=(1,1),\qquad v^{\prime}=(b^{-},b^{+}).$
We now treat the nontrivial case $a^{+}>a^{-}$. Define
$I_{+}:=\bigl\{i:y_{i}^{*}/x_{i}=a^{+}\bigr\},\qquad I_{-}:=\bigl\{i:y_{i}^{*}/x_{i}=a^{-}\bigr\},$
which form a partition of $\{1,\dots,d\}$. Set
$x_{\pm}:=\sum_{i\in I_{\pm}}x_{i},\qquad u_{\pm}:=\frac{1}{x_{\pm}}\sum_{i\in I_{\pm}}x_{i}u_{i},$
so that
$x_{\pm}u_{\pm}=\sum_{i\in I_{\pm}}x_{i}u_{i}.$
Set
$y_{\pm}:=a^{\pm}x_{\pm},\qquad v_{\pm}:=b^{\pm}u_{\pm},$
and define the two-dimensional vectors
$x^{\prime}:=(x_{-},x_{+}),\qquad y^{\prime}:=(y_{-},y_{+}),\qquad u^{\prime}:=(u_{-},u_{+}),\qquad v^{\prime}:=(v_{-},v_{+}).$
A direct computation using the above definitions gives
$x^{\prime}\cdot u^{\prime}=x_{-}u_{-}+x_{+}u_{+}=\sum_{i\in I_{-}}x_{i}u_{i}+\sum_{i\in I_{+}}x_{i}u_{i}=x\cdot u,$
and similarly
$x^{\prime}\cdot v^{\prime}=x\cdot v^{*},\qquad y^{\prime}\cdot u^{\prime}=y^{*}\cdot u,\qquad y^{\prime}\cdot v^{\prime}=y^{*}\cdot v^{*}.$
Hence
$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})=\frac{(x^{\prime}\cdot u^{\prime})(y^{\prime}\cdot v^{\prime})}{(x^{\prime}\cdot v^{\prime})(y^{\prime}\cdot u^{\prime})}=\frac{(x\cdot u)(y^{*}\cdot v^{*})}{(x\cdot v^{*})(y^{*}\cdot u)}=F_{d}(x,y^{*};u,v^{*}),$
which completes the proof.
Corollary 3.3 (Four-Point Collapse).
Let $d\geq 2$ and $x,y,u,v\in\mathbb{R}^{d}_{>0}$. Then there exist $x^{\prime},y^{\prime},u^{\prime},v^{\prime}\in\mathbb{R}^{2}_{>0}$ such that
$s^{\pm}(x^{\prime},y^{\prime})=s^{\pm}(x,y),\qquad s^{\pm}(u^{\prime},v^{\prime})=s^{\pm}(u,v),$
and
$F_{d}(x,y;u,v)\;\leq\;F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime}).$
Proof.
Apply lemma 3.1 to obtain $y^{*},v^{*}$ with
$F_{d}(x,y^{*};u,v^{*})\;\geq\;F_{d}(x,y;u,v),$
then lemma 3.2 to $(x,y^{*};u,v^{*})$ to obtain $x^{\prime},y^{\prime},u^{\prime},v^{\prime}\in\mathbb{R}^{2}_{>0}$ with
$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})=F_{d}(x,y^{*};u,v^{*}),$
which gives the claim. ∎
Theorem 3.4 (Envelope Inequality in Arbitrary Dimension).
Let $A\in\mathbb{R}^{d_{1}\times d_{2}}_{>0}$ and $B\in\mathbb{R}^{d_{2}\times d_{3}}_{>0}$, where $d_{1},d_{2},d_{3}\geq 2$. Then
$\boxed{\,R(AB)\;\leq\;\Phi\big(R(A),R(B)\big)\,}.$
Proof.
By definition, $R(AB)$ is the maximum distortion among all $2\times 2$ sub-blocks of $AB$. Choose rows $i,j$ of $A$ and columns $k,\ell$ of $B$ such that this maximum is attained.
Let
[…]
Then the corresponding $2\times 2$ block of $AB$ is
$\begin{pmatrix}x\cdot u&x\cdot v\\ y\cdot u&y\cdot v\end{pmatrix}$,
whose distortion is
$R(AB)\;=\;F_{d}(x,y;u,v).$
Set
$\alpha:=\operatorname{Dist}(x,y)\leq R(A),\qquad\beta:=\operatorname{Dist}(u,v)\leq R(B).$
By Corollary 3.3, there exist $x^{\prime},y^{\prime},u^{\prime},v^{\prime}\in\mathbb{R}_{>0}^{2}$ such that
$\operatorname{Dist}(x^{\prime},y^{\prime})=\operatorname{Dist}(x,y)=\alpha,\qquad\operatorname{Dist}(u^{\prime},v^{\prime})=\operatorname{Dist}(u,v)=\beta,$
and
$F_{d}(x,y;u,v)\;\leq\;F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime}).$
Let $A^{\prime}$ be the $2\times 2$ matrix with rows $x^{\prime}$ and $y^{\prime}$, and let $B^{\prime}$ be the $2\times 2$ matrix with columns $u^{\prime}$ and $v^{\prime}$. Then $R(A^{\prime})=\operatorname{Dist}(x,y)=\alpha$ and $R(B^{\prime})=\operatorname{Dist}(u,v)=\beta$. Then
$A^{\prime}B^{\prime}=\begin{pmatrix}x^{\prime}\!\cdot u^{\prime}&x^{\prime}\!\cdot v^{\prime}\\[2.0pt] y^{\prime}\!\cdot u^{\prime}&y^{\prime}\!\cdot v^{\prime}\end{pmatrix}$,
and therefore
$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})\;=\;R(A^{\prime}B^{\prime}).$
Combining the above steps and using that $\Phi$ is increasing in each variable gives
$R(AB)$ $=F_{d}(x,y;u,v)$
$\leq F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})$
$=R(A^{\prime}B^{\prime})$
$\leq\Phi(\alpha,\beta)$
$\leq\Phi(R(A),R(B)).$
This completes the proof.
4 Distortion of Products
Let $(A_{n})_{n\geq 1}$ be a sequence of positive matrices and define the partial products
$P_{n}:=A_{n}\cdots A_{1}.$
The matrices may have arbitrary dimensions, provided the products are well defined.
It is convenient to work with the square root of the distortion,
$S(A):=\sqrt{R(A)}.$
The envelope inequality then takes the simpler form
$S(AB)\;\leq\;\Psi\bigl(S(A),S(B)\bigr),\qquad\Psi(p,q):=\frac{1+pq}{p+q}.$
Since $\Psi$ is increasing in each variable, distortion bounds propagate along the product. If
$R(A_{n})\leq\alpha_{n},\qquad\text{i.e.}\qquad S(A_{n})\leq p_{n}:=\sqrt{\alpha_{n}},$
then
$S(P_{n})\leq q_{n},$
where $(q_{n})$ is defined recursively by
$q_{1}=p_{1},\qquad q_{n+1}=\Psi(p_{n+1},q_{n}).$
Finally,
$R(P_{n})\leq q_{n}^{2}.$
Uniformly bounded distortions.
Assume that all matrices satisfy the uniform bound
$R(A_{n})\leq\alpha,\qquad\text{i.e.}\qquad S(A_{n})\leq p:=\sqrt{\alpha}.$
Then the upper bounds $q_{n}$ satisfy
$q_{n+1}=\Psi(p,q_{n}),\qquad q_{1}=p.$
With $p$ fixed, this iterates a single Möbius map. Writing
$\kappa:=\frac{p-1}{p+1},$
its $n$th iterate can be expressed as
$q_{n}=1+\frac{2\,\kappa^{n}}{1-\kappa^{n}}.$
Consequently,
$R(P_{n})\leq q_{n}^{2}=1+\frac{4\,\kappa^{n}}{(1-\kappa^{n})^{2}}.$
For convenience, we record the following standalone theorem.
Theorem 4.1 (Finite-Product Distortion Bound).
Let $(A_{n})_{n\geq 1}$ be a sequence of positive matrices, possibly of varying dimensions, and assume that all products $P_{n}:=A_{n}\cdots A_{1}$ are well defined.
For any positive matrix $B>0$, define its distortion by
$R(B):=\max_{i,j,k,\ell}\frac{b_{ik}\,b_{j\ell}}{b_{i\ell}\,b_{jk}}.$
Assume that all factors satisfy the uniform bound
$R(A_{n})\leq\alpha\qquad\text{for all }n\geq 1.$
Set
$\kappa:=\frac{\sqrt{\alpha}-1}{\sqrt{\alpha}+1}.$
Then, for every $n\geq 1$,
$R(P_{n})\;\leq\;1\;+\;\frac{4\,\kappa^{\,n}}{(1-\kappa^{\,n})^{2}},$
and therefore
$R(P_{n})\longrightarrow 1\quad\text{at least exponentially fast as }n\to\infty.$
5 Comparison with the Birkhoff–Bushell
Contraction Theory
In classical Birkhoff–Bushell (BB) theory one studies distortion on a log scale via the Hilbert metric
$d_{H}(x,y)=\log\mathrm{Dist}(x,y),\qquad x,y\in\mathbb{R}^{n}_{>0}.$
If $B$ is an $n\times 2$ matrix with columns $x$ and $y$, and $A$ is any positive $n\times n$ matrix, then
$d_{H}(x,y)=\log R(B),\qquad d_{H}(Ax,Ay)=\log R(AB).$
In this log setting, our envelope inequality becomes
$d_{H}(Ax,Ay)\;\leq\;\Theta\!\left(d_{H}(x,y)\right),$
where
$\Theta(h)=2\log\!\left(\frac{1+p\,e^{h/2}}{\,p+e^{h/2}}\right),\qquad p=\sqrt{R(A)}.$
In the nontrivial case $p>1$, a direct computation shows that $\Theta$ is strictly increasing, strictly concave, and satisfies
$\Theta(0)=0,\qquad\Theta^{\prime}(0)=\frac{p-1}{p+1}=:\kappa(A).$
Thus $\Theta$ lies below its tangent line at the origin:
$\Theta(h)\;\leq\;\kappa(A)\,h,$
which yields the classical BB contraction:
$d_{H}(Ax,Ay)\;\leq\;\kappa(A)\,d_{H}(x,y).$
Note that
$\lim_{h\to\infty}\Theta(h)=\log R(A),$
so the envelope bound $\Theta(h)$ remains uniformly bounded, whereas the BB linear bound
$\kappa(A)\,h$
diverges. Thus both bounds agree to first order at $h=0$ (and therefore yield the same Lyapunov–gap asymptotics for distortion of high powers of $A$), but for large $h$ the BB estimate increasingly overestimates the worst–case contraction allowed by $R(A)$, while the envelope bound remains sharp in the non-asymptotic regime.
For illustration, Figure 2 compares the nonlinear envelope map $h\mapsto\Theta(h)$ with the classical BB line $k(A)h$, together with the saturation level $\log R(A)$.
[…]
Figure 2: Comparison of the envelope $\Theta(h)$ with the classical Birkhoff–Bushell contraction bound $\kappa(A)h$ for $R(A)=9$. The envelope contraction saturates at $\log R(A)$, while the BB line grows without bound, showing that the BB estimate increasingly overestimates the worst–case contraction allowed by $R(A)$ for large $h$.
Historical perspective.
Birkhoff’s fundamental insight was that positive operators act as contractions in the Hilbert metric, allowing the Frobenius–Perron alignment phenomenon to be deduced from the Banach fixed-point theorem. Bushell later refined this viewpoint by computing an explicit contraction constant. In most contraction theorems such constants are technically useful but algebraically unremarkable: they serve as quantitative bounds and carry little intrinsic meaning. Bushell’s constant is a striking exception. Its closed form,
$\kappa(A)=\frac{\sqrt{R(A)}-1}{\sqrt{R(A)}+1},$
is unexpectedly symmetric and rigid—far too elegant, in retrospect, to be merely a sharp Lipschitz coefficient. Moreover, the metric framework of the Hilbert metric plays no essential role in the classical proof: it is used only to invoke the Banach fixed-point theorem, not to explain the mechanism of alignment itself. This indicates that the contraction phenomenon behind the Frobenius–Perron theory is not inherently metric but fundamentally algebraic, encoded in the behaviour of ratios rather than in the geometry induced by the Hilbert metric.
This motivates our approach: the envelope inequality shows that worst-case distortion evolves under Möbius iterations with a common attracting fixed point at $1$. In this formulation, alignment arises from the explicit fixed-point dynamics of these maps rather than from a geometric contraction in the Hilbert metric.
6 Discussion and outlook
The envelope inequality is dimension-free, and the same collapse mechanism extends to a broad class of infinite matrices and positive operators on cones. The operator-theoretic generalizations will be developed elsewhere.
A second direction concerns other projective quantities. Since the collapse argument in Section 3 uses only the four-point ratios
$\frac{(x\!\cdot\!u)(y\!\cdot\!v)}{(x\!\cdot\!v)(y\!\cdot\!u)},$
it is natural to conjecture that analogous two-dimensional collapse mechanisms may apply to a broader class of projective functionals.
The envelope inequality also suggests applications in Markov chain theory. Classical asymptotic mixing bounds typically rely on Doeblin, Dobrushin, or Hilbert-metric methods, whereas the envelope provides an exact non-asymptotic contraction profile. This may offer an alternative route to quantitative bounds for finite-time mixing of positive stochastic matrices.
The recursion in Section 4 provides a deterministic model for worst-case distortion in products of positive matrices; its transfer-matrix form makes the evolution completely explicit. It may also serve as a foundation for combining our methods with probabilistic techniques to study non-asymptotic behavior of random products.
These directions lie beyond the scope of the present paper. Our purpose here is purely foundational: to isolate and prove the core envelope inequality in its simplest form.
Appendix A Calculus Tools
In this appendix we prove properties of the functions (for $\alpha,\beta\geq 1$)
$f_{\alpha,\beta}(t)=\frac{(1+t)(1+\alpha\beta\,t)}{(1+\alpha t)(1+\beta t)},\qquad t>0,$
and
$\Phi(\alpha,\beta)=\left(\frac{1+\sqrt{\alpha\beta}}{\sqrt{\alpha}+\sqrt{\beta}}\right)^{2}.$
Lemma A.1.
The function $f_{\alpha,\beta}$ has the following properties:
  1. 1.
    $f_{\alpha,\beta}(t)\geq 1$ for all $t>0$;
  2. 2.
    if $\alpha,\beta>1$, the unique global maximum is attained at
    $t^{*}=\frac{1}{\sqrt{\alpha\beta}},$
    and
    $f_{\alpha,\beta}(t)\leq f_{\alpha,\beta}(t^{*})=\Phi(\alpha,\beta);$
  3. 3.
    if $\alpha=1$ or $\beta=1$, then $f_{\alpha,\beta}(t)\equiv 1$.
Proof.
(1) Expanding $f_{\alpha,\beta}(t)\geq 1$ gives
$(1+t)(1+\alpha\beta t)\;\geq\;(1+\alpha t)(1+\beta t),$
which simplifies to
$(\alpha-1)(\beta-1)t\geq 0,$
valid for all $\alpha,\beta\geq 1$ and $t>0$.
(2) Set $\varphi(t)=\log f_{\alpha,\beta}(t)$. A straightforward algebraic simplification yields
$\varphi^{\prime}(t)=\,\frac{(\alpha-1)(\beta-1)\,(1-\alpha\beta t^{2})}{(1+t)(1+\alpha t)(1+\beta t)(1+\alpha\beta t)}.$
Thus $\varphi^{\prime}(t)=0$ exactly at
$t^{*}=\frac{1}{\sqrt{\alpha\beta}},$
and this is the unique global maximum. Substituting $t^{*}$ gives the stated value of $f_{\alpha,\beta}(t^{*})$.
(3) If $\alpha=1$ or $\beta=1$, the formula reduces to $f_{\alpha,\beta}(t)\equiv 1$. ∎
Lemma A.2.
For $\alpha,\beta\geq 1$, the envelope function $\Phi(\alpha,\beta)$ is increasing in each variable: for all $\alpha_{1}\leq\alpha_{2}$ and $\beta_{1}\leq\beta_{2}$,
$\Phi(\alpha_{1},\beta)\leq\Phi(\alpha_{2},\beta),\qquad\Phi(\alpha,\beta_{1})\leq\Phi(\alpha,\beta_{2}).$
Proof.
Write $p=\sqrt{\alpha}$ and $q=\sqrt{\beta}$, so $p,q\geq 1$, and define
$\psi(p,q)=\frac{1+pq}{p+q},$
so that $\Phi(\alpha,\beta)=\psi(p,q)^{2}$. For fixed $q\geq 1$ we compute
$\frac{\partial\psi}{\partial p}=\frac{q(p+q)-(1+pq)}{(p+q)^{2}}=\frac{q^{2}-1}{(p+q)^{2}}\;\geq\;0.$
Thus $\psi$ is increasing in $p$, and therefore so is $\Phi=\psi^{2}$. By symmetry the same argument applies to $q$. ∎

### v2

@@
 Abstract
-We introduce an elementary framework that captures the mechanism driving the alignment of rows and columns in products of positive matrices.
-All worst-case misalignment occurs already in dimension two, leading to an explicit collapse principle and a sharp nonlinear bound for finite products.
-The proof avoids Hilbert-metric and cone-theoretic techniques, relying instead on basic calculus.
-In the Hilbert metric, the classical Birkhoff–Bushell contraction captures only the linearized asymptotic regime, whereas our nonlinear envelope function gives the exact worst-case behavior for finite products.
+We study the deviation from proportionality of rows and columns in products of positive matrices.
+We prove a sharp, dimension-free bound showing that worst-case misalignment is already captured in dimension two and follows an explicit Möbius law, refining the classical Birkhoff–Bushell contraction theory.
 1 Introduction
-Products of positive matrices exhibit a pronounced alignment phenomenon: as the product grows, all rows and columns become nearly proportional.
-Classical methods approach this behavior through the Hilbert projective metric, probabilistic tools, or spectral theory.
-In this work we take a different route: we work directly with the distortion
-$R(A):=\max_{i,j,k,\ell}\frac{a_{ik}a_{j\ell}}{a_{i\ell}a_{jk}},\qquad A\in\mathbb{R}_{>0}^{d_{1}\times d_{2}},$
-which quantifies how far the rows or columns of $A$ are from being proportional.
-We have $R(A)=1$ in the perfectly aligned case; in general $R(A)\geq 1$ and increases with misalignment.
-Although $R(A)$ corresponds to the exponential of the projective diameter of $A$, we rely only on its elementary ratio form and use no projective-metric theory.
-Tracking how distortion changes under matrix multiplication gives an elementary, coordinate-level mechanism for alignment.
-The central observation is that worst–case distortion occurs already in dimension two; higher dimensions do not produce any new extremal behaviour.
-In dimension two, the explicit sharp distortion bound can then be computed by an elementary one–variable calculus optimization.
-A dimension collapse argument shows that exactly the same bound holds in arbitrary dimension.
-This yields the envelope inequality
-$R(AB)\leq\Phi(R(A),R(B)),$
-where
-$\Phi(\alpha,\beta)=\left(\frac{1+\sqrt{\alpha\beta}}{\sqrt{\alpha}+\sqrt{\beta}}\right)^{\!2}.$
-Roadmap.
-Section 2 introduces distortion and proves the envelope inequality in dimension 2.
-Section 3 extends the envelope inequality to arbitrary dimension via a collapse argument.
-Section 4 develops distortion for longer products of matrices, including explicit Frobenius–Perron bounds.
-Section 5 compares our result with the classical Birkhoff–Bushell theory.
-Section 6 highlights several directions for further work.
-Additional calculus lemmas used in the proofs are collected in the Appendix.
-2 Distortion Calculus in Dimension 22
+Products of positive matrices exhibit a pronounced alignment phenomenon: as the product grows, all columns (respectively, all rows) become nearly proportional.
+While alignment is well understood asymptotically, the finite-time law governing worst-case misalignment has not yet been isolated.
+In this paper, we develop an elementary framework based on the distortion functional
+$R(A):=\max_{\begin{subarray}{c}1\leq i,j\leq d_{1}\\ 1\leq k,\ell\leq d_{2}\end{subarray}}\frac{a_{ik}\,a_{j\ell}}{a_{i\ell}\,a_{jk}},\qquad A\in\mathbb{R}_{>0}^{d_{1}\times d_{2}}.$
+Our main result is a sharp envelope inequality controlling the distortion of matrix products,
+$R(AB)\leq\Phi(R(A),R(B)),\qquad\Phi(\alpha,\beta):=\left(\frac{1+\sqrt{\alpha\beta}}{\sqrt{\alpha}+\sqrt{\beta}}\right)^{2}.$
+A key structural observation is a rigidity phenomenon for projective distortion: extremal misalignment does not engage genuinely high-dimensional geometry, but collapses to a two-dimensional extremal configuration.
+While distortion is classically measured by the projective diameter $\log R(A)$ in the Hilbert metric, we show that $\sqrt{R(A)}$ is the natural scale on which extremal distortion evolves according to an explicit Möbius law.
+Organization of the paper.
+Section 2 establishes the envelope inequality in dimension 2.
+Section 3 extends the result to arbitrary dimension via a dimension-reduction argument.
+Section 4 derives explicit finite-time bounds for the distortion of matrix products.
+Section 5 compares the envelope bound with the classical Birkhoff–Bushell theory.
+2 Distortion Bounds in Dimension 22
 Definition 2.1.
@@
     $R(A):=\max\{F(A),\,1/F(A)\}.$
-Both $F(A)$ and $R(A)$ measure how far the two columns (equivalently, the two rows) of $A$ are from being proportional.
-Perfect proportionality is equivalent to $F(A)=R(A)=1$.
-The following properties are immediate from the definition:
-  • Scaling invariance. Multiplying any row or any column of $A$ by a positive scalar leaves both $F(A)$ and $R(A)$ unchanged.
-  • Effect of swapping. Swapping the two rows or the two columns of $A$ replaces $F(A)$ by $1/F(A)$ and therefore leaves $R(A)$ unchanged.
+Both $F(A)$ and $R(A)$ measure the deviation of the two columns (or rows) of $A$ from proportionality. Perfect proportionality is equivalent to $F(A)=R(A)=1$. The following properties are immediate.
+  • Scaling invariance. Multiplying any row or column of $A$ by a positive scalar leaves $F(A)$ and $R(A)$ unchanged.
+  • Effect of swapping. Swapping the two rows or the two columns replaces $F(A)$ by $1/F(A)$ and leaves $R(A)$ unchanged.
   • Orientation via determinant. For $2\times 2$ positive matrices,
@@
     $R(A)=1/F(A)\iff F(A)\leq 1\iff\det(A)\leq 0.$
-    In the perfectly aligned case, $\det(A)=0$ and both orientations coincide.
-Note that in our projective, scale-invariant setting, the area interpretation of the determinant is irrelevant; only the sign information matters.
-Our goal is to obtain a sharp explicit upper bound on the distortion
-$R(AB)$
-of the product of two positive $2\times 2$ matrices. The analysis is greatly simplified by a few distortion-preserving normalizations.
-We first ensure positive orientation:
+    In the proportional case $\det(A)=0$, both orientations coincide.
+Our goal is to obtain a sharp explicit upper bound on $R(AB)$ for the product of two positive $2\times 2$ matrices. The analysis is simplified by distortion-preserving normalizations.
+We first ensure positive orientation,
 $F(A)\geq 1,\qquad F(B)\geq 1,$
 so that $F(A)=R(A)$ and $F(B)=R(B)$.
-This can always be achieved by swapping the rows of $A$ and/or the columns of $B$; such swaps do not affect the distortion of $A$, $B$, or $AB$.
-After ensuring positive orientation of $A$ and $B$, we have $\det(A)\geq 0$, $\det(B)\geq 0$, and so $\det(AB)=\det(A)\det(B)\geq 0$.
-Then $F(AB)\geq 1$ and $AB$ is also positively oriented: $F(AB)=R(AB)$.
-Next, we are free to multiply the rows of $A$ and the columns of $B$ by any positive scalars. Such scalings affect neither the orientation nor the distortion of $A$, $B$, or $AB$.
-After these normalizations, we may assume, without loss of generality, that $A$ and $B$ are in the canonical form
+This can be achieved by swapping rows of $A$ and/or columns of $B$, which does not affect the distortions of $A$, $B$, or $AB$.
+Then $\det(A),\det(B)\geq 0$, and hence $\det(AB)\geq 0$, so that $F(AB)=R(AB)$.
+We may also multiply the rows of $A$ and the columns of $B$ by arbitrary positive scalars. Such scalings affect neither the orientation nor the distortion of $A$, $B$, or $AB$.
+After these normalizations, we may assume without loss of generality that
 $A=\begin{pmatrix}1&u\\[5.69054pt] 1&\alpha u\end{pmatrix},\qquad B=\begin{pmatrix}1&1\\[5.69054pt] v&\beta v\end{pmatrix}$,
@@
 $\alpha=R(A)\geq 1,\qquad\beta=R(B)\geq 1.$
-A direct computation gives
-$AB=\begin{pmatrix}1+uv&1+\beta\,uv\\[5.69054pt] 1+\alpha\,uv&1+\alpha\beta\,uv\end{pmatrix}=\begin{pmatrix}1+t&1+\beta t\\[5.69054pt] 1+\alpha t&1+\alpha\beta t\end{pmatrix}$,
+Then
+$AB=\begin{pmatrix}1+t&1+\beta t\\[5.69054pt] 1+\alpha t&1+\alpha\beta t\end{pmatrix},\qquad t=uv>0.$
+Accordingly,
+$R(AB)=f_{\alpha,\beta}(t),\qquad f_{\alpha,\beta}(t):=\frac{(1+t)(1+\alpha\beta t)}{(1+\beta t)(1+\alpha t)},\quad t>0.$
+The task of bounding $R(AB)$ therefore reduces to maximizing the one-variable function $f_{\alpha,\beta}(t)$. Differentiating its logarithm yields
+$\frac{d}{dt}\log f_{\alpha,\beta}(t)=\frac{(\alpha-1)(\beta-1)\,(1-\alpha\beta t^{2})}{(1+t)(1+\alpha t)(1+\beta t)(1+\alpha\beta t)}.$
+Thus $f_{\alpha,\beta}$ is maximized at
+$t^{*}:=\frac{1}{\sqrt{\alpha\beta}},$
+with
+$f_{\alpha,\beta}(t^{*})=\Phi(\alpha,\beta),$
 where
-$t=uv>0.$
-Then
-$R(AB)=f_{\alpha,\beta}(uv),$
-where
-$f_{\alpha,\beta}(t):=\frac{(1+t)(1+\alpha\beta t)}{(1+\beta t)(1+\alpha t)}\qquad t>0.$
-Thus the problem of finding a sharp upper bound for $R(AB)$ in terms of $R(A)$ and $R(B)$ reduces to a simple calculus exercise: it suffices to maximize the function $f_{\alpha,\beta}(t)$ for $t>0$.
-The details are carried out in Lemma A.1 and the result is
-$\max_{t>0}f_{\alpha,\beta}(t)=\Phi(\alpha,\beta),$
-where
-$\Phi(\alpha,\beta)=\frac{\bigl(1+\sqrt{\alpha\beta}\bigr)^{2}}{\bigl(\sqrt{\alpha}+\sqrt{\beta}\bigr)^{2}}.$
-We refer to $\Phi$ as the envelope function.
-Thus we obtain
-$R(AB)=f_{\alpha,\beta}(uv)\leq\Phi(\alpha,\beta),$
-which proves the following bound:
+$\Phi(\alpha,\beta):=\left(\frac{1+\sqrt{\alpha\beta}}{\sqrt{\alpha}+\sqrt{\beta}}\right)^{2}.$
+We refer to $\Phi(\alpha,\beta)$ as the envelope function. Hence
+$R(AB)\leq\Phi(\alpha,\beta),$
+and we obtain the following result.
 Theorem 2.2 (Envelope Inequality in Dimension $2$).
 Let $A$ and $B$ be strictly positive $2\times 2$ matrices. Then
-$\boxed{R(AB)\leq\Phi\bigl(R(A),R(B)\bigr)=\left(\frac{1+\sqrt{R(A)R(B)}}{\sqrt{R(A)}+\sqrt{R(B)}}\right)^{\!2}}.$
-Moreover, the envelope function $\Phi(\alpha,\beta)$ is increasing in each variable (Lemma A.2). Therefore, from Theorem 2.2 we obtain:
+$R(AB)\leq\Phi\bigl(R(A),R(B)\bigr)=\left(\frac{1+\sqrt{R(A)R(B)}}{\sqrt{R(A)}+\sqrt{R(B)}}\right)^{\!2}.$
+The envelope function $\Phi$ is increasing in each variable, yielding the following corollary.
 Corollary 2.3.
-Let $A$ and $B$ be strictly positive $2\times 2$ matrices, and suppose
-$R(A)\leq\alpha,\qquad R(B)\leq\beta.$
-Then
+Let $A$ and $B$ be strictly positive $2\times 2$ matrices with $R(A)\leq\alpha$ and $R(B)\leq\beta$. Then
 $R(AB)\leq\Phi(\alpha,\beta).$
-We note that the envelope inequality is not merely an upper bound; the maximal distortion is attainable.
+The envelope inequality is sharp.
 Proposition 2.4 (Two–Sided Sharpness).
@@
   1. (i)
-    Given any $A>0$ with $R(A)=\alpha$, one can choose $B>0$ with $R(B)=\beta$ such that
-    $R(AB)=\Phi(\alpha,\beta).$
+    For any $A>0$ with $R(A)=\alpha$, there exists $B>0$ with $R(B)=\beta$ such that $R(AB)=\Phi(\alpha,\beta)$.
   2. (ii)
-    Given any $B>0$ with $R(B)=\beta$, one can choose $A>0$ with $R(A)=\alpha$ such that
-    $R(AB)=\Phi(\alpha,\beta).$
+    For any $B>0$ with $R(B)=\beta$, there exists $A>0$ with $R(A)=\alpha$ such that $R(AB)=\Phi(\alpha,\beta)$.
 Proof.
-We first prove (i). By Lemma A.1, the function $f_{\alpha,\beta}(t)$ attains its maximum $\Phi(\alpha,\beta)$ at
-$t^{*}=\frac{1}{\sqrt{\alpha\beta}}.$
-Every product of the form
-$\begin{pmatrix}1&u\\[5.69054pt] 1&\alpha u\end{pmatrix}\begin{pmatrix}1&1\\[5.69054pt] v&\beta v\end{pmatrix},\qquad uv=t^{*},$
+We first prove (i). Every product of the form
+$\begin{pmatrix}1&u\\ 1&\alpha u\end{pmatrix}\begin{pmatrix}1&1\\ v&\beta v\end{pmatrix},\qquad uv=t^{*},$
 has distortion $\Phi(\alpha,\beta)$.
-Let $A>0$ be given with $R(A)=\alpha$. By rescaling the rows of $A$ by positive factors (which does not change $R(A)$ nor $R(AB)$), we may assume that
-$A=\begin{pmatrix}1&x\\[5.69054pt] 1&y\end{pmatrix},\qquad x,y>0.$
-If $A$ is positively oriented, i.e. $y/x=\alpha$, we take
-$B=\begin{pmatrix}1&1\\[5.69054pt] v&\beta v\end{pmatrix},\qquad v=\frac{t^{*}}{x}.$
-If $A$ is negatively oriented, i.e. $x/y=\alpha$, we take
-$B=\begin{pmatrix}1&1\\[5.69054pt] \beta v&v\end{pmatrix},\qquad v=\frac{t^{*}}{y}.$
-In both cases one checks directly that
-$R(AB)=\Phi(\alpha,\beta).$
-This proves (i).
-For (ii), use that $R(M)=R(M^{\mathsf{T}})$ and $\Phi(\alpha,\beta)=\Phi(\beta,\alpha)$.
-Given $B$ with $R(B)=\beta$, apply (i) to $B^{\mathsf{T}}$ and $\alpha$: there exists $C>0$ with $R(C)=\alpha$ such that
+Let $A>0$ satisfy $R(A)=\alpha$. By rescaling rows, we may write
+$A=\begin{pmatrix}1&x\\ 1&y\end{pmatrix},\qquad x,y>0.$
+If $A$ is positively oriented, so that $y/x=\alpha$, we take
+$B=\begin{pmatrix}1&1\\ v&\beta v\end{pmatrix},\qquad v=\frac{t^{*}}{x}.$
+If $A$ is negatively oriented, so that $x/y=\alpha$, we take
+$B=\begin{pmatrix}1&1\\ \beta v&v\end{pmatrix},\qquad v=\frac{t^{*}}{y}.$
+In both cases, $R(AB)=\Phi(\alpha,\beta)$, proving (i).
+For (ii), we use the symmetry under transposition. Note that
+$R(M)=R(M^{\mathsf{T}})\qquad\text{and}\qquad\Phi(\alpha,\beta)=\Phi(\beta,\alpha).$
+Given $B>0$ with $R(B)=\beta$, apply part (i) to $B^{\mathsf{T}}$. There exists $C>0$ with $R(C)=\alpha$ such that
 $R(B^{\mathsf{T}}C)=\Phi(\beta,\alpha).$
@@
 For all $\alpha,\beta\geq 1$,
-$\boxed{\Phi(\alpha,\beta)=\max\bigl\{\,R(AB):A,B>0,\ R(A)\leq\alpha,\ R(B)\leq\beta\,\bigr\}}.$
-3 Reduction to Two Dimensions in the General Case
-The goal of this section is to extend the envelope inequality to arbitrary dimension.
+$\Phi(\alpha,\beta)=\max\bigl\{R(AB):A,B>0,\ R(A)\leq\alpha,\ R(B)\leq\beta\bigr\}.$
+3 Reduction to Two Dimensions
+In this section we extend the notion of distortion to arbitrary dimension and prove the envelope inequality via a reduction to two dimensions.
 For $x,y\in\mathbb{R}_{>0}^{d}$, define
-$s^{+}(x,y)=\max_{i}\frac{y_{i}}{x_{i}},\qquad s^{-}(x,y)=\min_{i}\frac{y_{i}}{x_{i}},$
-and the distortion
-$\operatorname{Dist}(x,y)=\frac{s^{+}(x,y)}{s^{-}(x,y)}.$
-For perfectly aligned vectors we have $s^{+}(x,y)=s^{-}(x,y)$, so $\operatorname{Dist}(x,y)=1$.
-When the ratios $y_{i}/x_{i}$ are more scattered, the distortion increases.
-Figure 1 illustrates $s^{+}(x,y)$ and $s^{-}(x,y)$ as maximal and minimal slopes of the points $(x_{i},y_{i})$ in $\mathbb{R}_{>0}^{2}$.
-Figure 1: The pair $(x,y)$ represented as a cloud of points between the extremal rays of slopes $s^{+}(x,y)$ and $s^{-}(x,y)$ in $\mathbb{R}_{>0}^{2}$.
+$s^{+}(x,y):=\max_{i}\frac{y_{i}}{x_{i}},\qquad s^{-}(x,y):=\min_{i}\frac{y_{i}}{x_{i}}.$
+Figure 1 illustrates $s^{\pm}(x,y)$ as the extremal slopes of the points $(x_{i},y_{i})$ in $\mathbb{R}_{>0}^{2}$.
+Figure 1: Geometric representation of $(x,y)$ as points in $\mathbb{R}_{>0}^{2}$.
+The misalignment of $x$ and $y$ is measured by the distortion
+$\operatorname{Dist}(x,y):=\frac{s^{+}(x,y)}{s^{-}(x,y)}.$
+Equivalently,
+$\operatorname{Dist}(x,y)=\max_{i,j}\frac{y_{i}x_{j}}{y_{j}x_{i}}.$
 We extend the definition of distortion to positive matrices of size $d_{1}\times d_{2}$ by
-$R(A)=\max_{\begin{subarray}{c}1\leq i,j\leq d_{1}\\ 1\leq k,\ell\leq d_{2}\end{subarray}}\frac{a_{ik}\,a_{j\ell}}{a_{i\ell}\,a_{jk}}.$
-Thus $R(A)$ is the maximum of the distortions of all $2\times 2$ submatrices of $A$, obtained by choosing any two rows and any two columns.
-Equivalently, it is the maximum distortion between any two columns (or two rows) of $A$:
+$R(A):=\max_{\begin{subarray}{c}1\leq i,j\leq d_{1}\\ 1\leq k,\ell\leq d_{2}\end{subarray}}\frac{a_{ik}\,a_{j\ell}}{a_{i\ell}\,a_{jk}}.$
+Thus $R(A)$ is the maximum distortion of all $2\times 2$ submatrices of $A$. Equivalently,
 $R(A)=\max_{k,\ell}\operatorname{Dist}\bigl(A_{*k},A_{*\ell}\bigr)=\max_{i,j}\operatorname{Dist}\bigl(A_{i*},A_{j*}\bigr).$
+The following functional encodes the oriented distortion of the $2\times 2$ block of a matrix product, obtained by pairing rows $x,y$ of the left factor with columns $u,v$ of the right factor.
 Define $F_{d}:(\mathbb{R}^{d}_{>0})^{4}\to\mathbb{R}_{>0}$ by
-$F_{d}(x,y;u,v):=\frac{(x\cdot u)(y\cdot v)}{(x\cdot v)(y\cdot u)}.$
-Here $x\cdot u$ denotes the standard Euclidean dot product.
+$F_{d}(x,y;u,v):=\frac{(x\cdot u)(y\cdot v)}{(x\cdot v)(y\cdot u)},$
+where $\cdot$ denotes the standard Euclidean dot product.
 Lemma 3.1 (coordinatewise extremal ratio pushing).
-Let $d\geq 2$ and let $x,y,u,v\in\mathbb{R}^{d}_{>0}$. Define
+Let $x,y,u,v\in\mathbb{R}^{d}_{>0}$ and set
 $a^{\pm}:=s^{\pm}(x,y),\qquad b^{\pm}:=s^{\pm}(u,v).$
@@
 $s^{\pm}(x,y^{*})=s^{\pm}(x,y),\qquad s^{\pm}(u,v^{*})=s^{\pm}(u,v),$
-and for all $i$,
+and for all $1\leq i\leq d$,
 $\frac{y_{i}^{*}}{x_{i}}\in\{a^{-},a^{+}\},\qquad\frac{v_{i}^{*}}{u_{i}}\in\{b^{-},b^{+}\}.$
 Proof.
-We will first construct $y^{*}$ and then use a similar procedure to obtain $v^{*}$.
-The strategy is to adjust the coordinates of $y$ one at a time, always modifying them in a way that does not decrease the value of $F_{d}(x,y;u,v)$, and while keeping the extremal slopes $s^{\pm}(x,y)$ unchanged.
-To make this precise, partition the index set $\{1,\dots,d\}$ into
-$I_{-}:=\{\,i:y_{i}/x_{i}=a^{-}\,\},\qquad I_{+}:=\{\,i:y_{i}/x_{i}=a^{+}\,\},$
+We first construct $y^{*}$. The strategy is to use monotonicity to adjust the coordinates of $y$ one at a time without decreasing the value of $F_{d}$.
+Define
+$f(z_{1},\dots,z_{d})\;:=\;F_{d}(x,z;u,v),\qquad z=(z_{1},\dots,z_{d})\in\mathbb{R}_{>0}^{d}.$
+For $i\in\{1,\dots,d\}$, write $z_{\neq i}:=(z_{1},\dots,z_{i-1},z_{i+1},\dots,z_{d})$ and
+$f_{i}(t\mid z_{\neq i})\;:=\;f(z_{1},\dots,z_{i-1},t,z_{i+1},\dots,z_{d}),\qquad t>0.$
+For fixed $z_{\neq i}$, the function $t\mapsto f_{i}(t\mid z_{\neq i})$ is fractional linear with strictly positive denominator, and hence is either increasing, decreasing, or constant.
+Set
+$B:=\{\,z\in\mathbb{R}_{>0}^{d}:a^{-}x_{i}\leq z_{i}\leq a^{+}x_{i}\text{ for all }i\,\}.$
+For each $i\in\{1,\dots,d\}$, define the coordinate update map $E_{i}$ on $B$ by
+[…]
+Then
+$E_{i}(z)\in\{a^{-}x_{i},a^{+}x_{i}\}\quad\text{for all }z\in B.$
+Define the associated vector update operator $\mathcal{E}_{i}:B\to B$ by
+$(\mathcal{E}_{i}z)_{j}:=\begin{cases}z_{j},&j\neq i,\\ E_{i}(z),&j=i.\end{cases}$
+By construction,
+[…]
+Define
+$y^{*}\;:=\;(\mathcal{E}_{d}\circ\cdots\circ\mathcal{E}_{1})(y).$
+Then
+$f(y^{*})\;\geq\;f(y),\qquad s^{\pm}(x,y^{*})\;=\;s^{\pm}(x,y),\qquad\frac{y_{i}^{*}}{x_{i}}\in\{a^{-},a^{+}\}\ \ \text{for all }i.$
+We next construct $v^{*}$ in the same manner. Define
+$g(w_{1},\dots,w_{d})\;:=\;F_{d}(x,y^{*};u,w),\qquad w=(w_{1},\dots,w_{d})\in\mathbb{R}_{>0}^{d},$
 and
-$I_{0}:=\{1,\dots,d\}\setminus(I_{-}\cup I_{+}).$
-We will never alter the coordinates in $I_{-}$ or $I_{+}$, so that the extremal slopes remain fixed at all times. Thus we immediately set
-$y_{i}^{*}:=y_{i}\qquad\text{for all }i\in I_{-}\cup I_{+}.$
-The coordinatewise replacement procedure is applied only to indices in $I_{0}$.
-We begin with a simple coordinate monotonicity observation. Fix $x,u,v$ and all coordinates of $y$ except $y_{i}$. Then
-$f(y_{i}):=F_{d}(x,y;u,v)$
-is a fractional linear function of $y_{i}$ with strictly positive denominator, and is therefore monotone (either increasing or decreasing) in $y_{i}$. If $f$ is increasing in $y_{i}$, we set
-$y_{i}^{*}:=a^{+}\,x_{i},$
-while if $f$ is decreasing in $y_{i}$, we set
-$y_{i}^{*}:=a^{-}\,x_{i}.$
+$g_{i}(t\mid w_{\neq i})\;:=\;g(w_{1},\dots,w_{i-1},t,w_{i+1},\dots,w_{d}),\qquad t>0.$
+For fixed $w_{\neq i}$, the function $t\mapsto g_{i}(t\mid w_{\neq i})$ is fractional linear with strictly positive denominator.
+Repeating the coordinate update argument in this setting produces $v^{*}\in\mathbb{R}_{>0}^{d}$ such that
+$g(v^{*})\;\geq\;g(v),\qquad s^{\pm}(u,v^{*})\;=\;s^{\pm}(u,v),\qquad\frac{v_{i}^{*}}{u_{i}}\in\{b^{-},b^{+}\}\ \ \text{for all }i.$
+Finally, by construction
+$F_{d}(x,y^{*};u,v^{*})$ $=g(v^{*})$
+$\geq g(v)$
+$=F_{d}(x,y^{*};u,v)$
+$=f(y^{*})$
+$\geq f(y)$
+$=F_{d}(x,y;u,v),$
+which completes the proof. ∎
+The following lemma shows that extremal configurations collapse to dimension two.
+Lemma 3.2 (Two-dimensional collapse).
+Let $x,y,u,v\in\mathbb{R}_{>0}^{d}$ and set
+$a^{\pm}:=s^{\pm}(x,y),\qquad b^{\pm}:=s^{\pm}(u,v).$
+Assume that
+$\frac{y_{i}}{x_{i}}\in\{a^{-},a^{+}\}\quad\text{and}\quad\frac{v_{i}}{u_{i}}\in\{b^{-},b^{+}\}\qquad\text{for all }1\leq i\leq d.$
+Then there exist vectors $x^{\prime},y^{\prime},u^{\prime},v^{\prime}\in\mathbb{R}_{>0}^{2}$ such that
+$s^{\pm}(x^{\prime},y^{\prime})=a^{\pm},\qquad s^{\pm}(u^{\prime},v^{\prime})=b^{\pm},$
+and
+$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})\;\geq\;F_{d}(x,y;u,v).$
+Proof.
+If $a^{-}=a^{+}=a$, then $y=ax$ and $F_{d}(x,y;u,v)=1$.
+In this case one may take $x^{\prime}=(1,1)$, $y^{\prime}=(a,a)$, $u^{\prime}=(1,1)$ and $v^{\prime}=(b^{-},b^{+})$, so that $s^{\pm}(x^{\prime},y^{\prime})=a$, $s^{\pm}(u^{\prime},v^{\prime})=b^{\pm}$, and $F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})=1$.
+The case $b^{-}=b^{+}$ is similar and is handled analogously.
+In either case the conclusion of the lemma holds.
+We therefore assume from now on that $a^{-}<a^{+}$ and $b^{-}<b^{+}$.
+Define the index sets
+$I_{-}:=\{\,i:\frac{y_{i}}{x_{i}}=a^{-}\,\},\qquad I_{+}:=\{\,i:\frac{y_{i}}{x_{i}}=a^{+}\,\},$
+which form a partition of $\{1,\dots,d\}$ and are both nonempty by assumption.
+Step 1: Two–dimensional reduction preserving $F_{d}$.
+Define
+$x^{-}:=\sum_{i\in I_{-}}x_{i},\qquad x^{+}:=\sum_{i\in I_{+}}x_{i},\qquad x^{\prime}:=(x^{-},x^{+})\in\mathbb{R}_{>0}^{2},$
+and
+$u^{-}:=\frac{1}{x^{-}}\sum_{i\in I_{-}}x_{i}u_{i},\qquad u^{+}:=\frac{1}{x^{+}}\sum_{i\in I_{+}}x_{i}u_{i},\qquad u^{\prime}:=(u^{-},u^{+})\in\mathbb{R}_{>0}^{2}.$
+Set
+$y^{\prime}:=(a^{-}x^{-},\,a^{+}x^{+})\in\mathbb{R}_{>0}^{2}.$
+By construction, $s^{\pm}(x^{\prime},y^{\prime})=a^{\pm}$. A direct computation shows that
+$x\cdot u=x^{\prime}\cdot u^{\prime},\qquad y\cdot u=y^{\prime}\cdot u^{\prime}.$
+Define an auxiliary vector $\tilde{v}\in\mathbb{R}_{>0}^{2}$ by
+[…]
 Then
-$F_{d}(x,y^{*};u,v)\geq F_{d}(x,y;u,v),\qquad s^{\pm}(x,y^{*})=s^{\pm}(x,y),$
+$x\cdot v=x^{\prime}\cdot\tilde{v},\qquad y\cdot v=y^{\prime}\cdot\tilde{v}$,
+and hence
+[…]
+Let
+$\tilde{b}^{-}:=\frac{\tilde{v}^{-}}{u^{-}},\qquad\tilde{b}^{+}:=\frac{\tilde{v}^{+}}{u^{+}}.$
+Then
+$\tilde{b}^{-}=\frac{1}{x^{-}u^{-}}\sum_{i\in I_{-}}x_{i}u_{i}\,\frac{v_{i}}{u_{i}},\qquad\tilde{b}^{+}=\frac{1}{x^{+}u^{+}}\sum_{i\in I_{+}}x_{i}u_{i}\,\frac{v_{i}}{u_{i}}.$
+Since $x^{-}u^{-}=\sum_{i\in I_{-}}x_{i}u_{i}$ and $\frac{v_{i}}{u_{i}}\in\{b^{-},b^{+}\}$, we conclude that $\tilde{b}^{-}$ is a weighted average of $b^{-}$ and $b^{+}$, and hence $\tilde{b}^{-}\in[b^{-},b^{+}]$.
+The same argument applies to $\tilde{b}^{+}$ and yields $\tilde{b}^{+}\in[b^{-},b^{+}]$.
+Note that in general $\tilde{b}^{\pm}\neq b^{\pm}$ so we cannot simply take $v^{\prime}:=\tilde{v}$ and conclude the proof. An extra step is needed.
+Step 2: Monotonicity argument.
+For $b_{1},b_{2}>0$ define
+$f(b_{1},b_{2}):=F_{2}\!\left(x^{\prime},y^{\prime};u^{\prime},(b_{1}u^{-},\,b_{2}u^{+})\right).$
+Then
+$f(b_{1},b_{2})=\frac{x^{\prime}\!\cdot u^{\prime}}{y^{\prime}\!\cdot u^{\prime}}\,\frac{Ab_{1}+Bb_{2}}{Cb_{1}+Db_{2}},$
+where
+$A:=a^{-}\,x^{-}u^{-},\qquad B:=a^{+}\,x^{+}u^{+},\qquad C:=x^{-}u^{-},\qquad D:=x^{+}u^{+}.$
+Note that
+$AD-BC=x^{-}u^{-}\,x^{+}u^{+}\,(a^{-}-a^{+})\;<\;0,$
+since $a^{-}<a^{+}$. Then
+$\frac{\partial f}{\partial b_{1}}(b_{1},b_{2})=\frac{x^{\prime}\!\cdot u^{\prime}}{y^{\prime}\!\cdot u^{\prime}}\,\frac{(AD-BC)\,b_{2}}{(Cb_{1}+Db_{2})^{2}}\;<\;0,$
+$\frac{\partial f}{\partial b_{2}}(b_{1},b_{2})=\frac{x^{\prime}\!\cdot u^{\prime}}{y^{\prime}\!\cdot u^{\prime}}\,\frac{(BC-AD)\,b_{1}}{(Cb_{1}+Db_{2})^{2}}\;>\;0,$
+for all $b_{1},b_{2}>0$. It follows that $f(b_{1},b_{2})$ is decreasing in $b_{1}$ and increasing in $b_{2}$.
+Define
+$v^{\prime}:=(b^{-}u^{-},\,b^{+}u^{+})\in\mathbb{R}_{>0}^{2},$
+so that $s^{\pm}(u^{\prime},v^{\prime})=b^{\pm}$. By construction,
+$f(b^{-},b^{+})=F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})$
 and
-$\frac{y_{i}^{*}}{x_{i}}\in\{\,a^{-},\,a^{+}\,\}.$
-The key point is that replacing $(x_{i},y_{i})$ by $(x_{i},y_{i}^{*})$ does not decrease $F_{d}$, does not change $s^{\pm}(x,y)$, and forces the ratio $y_{i}^{*}/x_{i}$ to be extremal.
-The frozen extremal indices in $I_{-}\cup I_{+}$ ensure that the original extremal ratios are preserved throughout.
-We now apply this replacement successively to all $i\in I_{0}$, always leaving previously modified coordinates unchanged. This yields a vector $y^{*}$ such that
-$s^{\pm}(x,y^{*})=s^{\pm}(x,y),\qquad F_{d}(x,y^{*};u,v)\geq F_{d}(x,y;u,v),$
-and every ratio $y_{i}^{*}/x_{i}$ is extremal:
-$\frac{y_{i}^{*}}{x_{i}}\in\{\,a^{-},\,a^{+}\,\}.$
-With $y^{*}$ fixed, we now apply exactly the same coordinatewise extremal-pushing procedure to the pair $(u,v)$.
-We freeze the indices where the ratios $v_{i}/u_{i}$ already attain the extremal values $b^{\pm}$, and modify only the remaining coordinates exactly as above: for each active coordinate we replace $v_{i}$ by either $b^{-}u_{i}$ or $b^{+}u_{i}$, according to the monotonicity of $F_{d}(x,y^{*};u,v)$ in $v_{i}$.
-This produces a vector $v^{*}$ such that
-$s^{\pm}(u,v^{*})=s^{\pm}(u,v),\qquad F_{d}(x,y^{*};u,v^{*})\geq F_{d}(x,y^{*};u,v),$
-and each ratio $v_{i}^{*}/u_{i}$ is extremal:
-$\frac{v_{i}^{*}}{u_{i}}\in\{\,b^{-},\,b^{+}\,\}.$
-This completes the proof. ∎
-Lemma 3.2 ($2$-point support collapse).
-Assume that the coordinate ratios of $y^{*}/x$ take only the two extremal values $a^{\pm}$, and the coordinate ratios of $v^{*}/u$ take only the two extremal values $b^{\pm}$, where
-$a^{\pm}=s^{\pm}(x,y^{*}),\qquad b^{\pm}=s^{\pm}(u,v^{*}).$
-Then there exist vectors $x^{\prime},y^{\prime},u^{\prime},v^{\prime}\in\mathbb{R}^{2}_{>0}$ such that
-$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})\;=\;F_{d}(x,y^{*};u,v^{*}),$
-and
-$s^{\pm}(x^{\prime},y^{\prime})=a^{\pm},\qquad s^{\pm}(u^{\prime},v^{\prime})=b^{\pm}.$
-Proof.
-If $a^{+}=a^{-}=a$, then $y^{*}=ax$ and hence $F_{d}(x,y^{*};u,v^{*})=1$. The lemma then holds, for example, with
-$x^{\prime}=(1,1),\qquad y^{\prime}=(a,a),\qquad u^{\prime}=(1,1),\qquad v^{\prime}=(b^{-},b^{+}).$
-We now treat the nontrivial case $a^{+}>a^{-}$. Define
-$I_{+}:=\bigl\{i:y_{i}^{*}/x_{i}=a^{+}\bigr\},\qquad I_{-}:=\bigl\{i:y_{i}^{*}/x_{i}=a^{-}\bigr\},$
-which form a partition of $\{1,\dots,d\}$. Set
-$x_{\pm}:=\sum_{i\in I_{\pm}}x_{i},\qquad u_{\pm}:=\frac{1}{x_{\pm}}\sum_{i\in I_{\pm}}x_{i}u_{i},$
-so that
-$x_{\pm}u_{\pm}=\sum_{i\in I_{\pm}}x_{i}u_{i}.$
-Set
-$y_{\pm}:=a^{\pm}x_{\pm},\qquad v_{\pm}:=b^{\pm}u_{\pm},$
-and define the two-dimensional vectors
-$x^{\prime}:=(x_{-},x_{+}),\qquad y^{\prime}:=(y_{-},y_{+}),\qquad u^{\prime}:=(u_{-},u_{+}),\qquad v^{\prime}:=(v_{-},v_{+}).$
-A direct computation using the above definitions gives
-$x^{\prime}\cdot u^{\prime}=x_{-}u_{-}+x_{+}u_{+}=\sum_{i\in I_{-}}x_{i}u_{i}+\sum_{i\in I_{+}}x_{i}u_{i}=x\cdot u,$
-and similarly
-$x^{\prime}\cdot v^{\prime}=x\cdot v^{*},\qquad y^{\prime}\cdot u^{\prime}=y^{*}\cdot u,\qquad y^{\prime}\cdot v^{\prime}=y^{*}\cdot v^{*}.$
-Hence
-$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})=\frac{(x^{\prime}\cdot u^{\prime})(y^{\prime}\cdot v^{\prime})}{(x^{\prime}\cdot v^{\prime})(y^{\prime}\cdot u^{\prime})}=\frac{(x\cdot u)(y^{*}\cdot v^{*})}{(x\cdot v^{*})(y^{*}\cdot u)}=F_{d}(x,y^{*};u,v^{*}),$
-which completes the proof.
+[…]
+Since $\tilde{b}^{\pm}\in[b^{-},b^{+}]$ and $f$ is decreasing in its first argument and increasing in its second, we conclude that
+$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})=f(b^{-},b^{+})\;\geq\;f(\tilde{b}^{-},\tilde{b}^{+})=F_{2}(x^{\prime},y^{\prime};u^{\prime},\tilde{v})=F_{d}(x,y;u,v).$
+This completes the proof.
 Corollary 3.3 (Four-Point Collapse).
@@
 and
-$F_{d}(x,y;u,v)\;\leq\;F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime}).$
+$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})\geq F_{d}(x,y;u,v).$
 Proof.
-Apply lemma 3.1 to obtain $y^{*},v^{*}$ with
-$F_{d}(x,y^{*};u,v^{*})\;\geq\;F_{d}(x,y;u,v),$
-then lemma 3.2 to $(x,y^{*};u,v^{*})$ to obtain $x^{\prime},y^{\prime},u^{\prime},v^{\prime}\in\mathbb{R}^{2}_{>0}$ with
-$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})=F_{d}(x,y^{*};u,v^{*}),$
-which gives the claim. ∎
+Apply Lemma 3.1 to obtain $y^{*},v^{*}$ with $F_{d}(x,y^{*};u,v^{*})\geq F_{d}(x,y;u,v)$.
+Then apply Lemma 3.2 to $x,y^{*},u$ and $v^{*}$ and obtain $x^{\prime},y^{\prime},u^{\prime},v^{\prime}\in\mathbb{R}^{2}_{>0}$ with $F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})\geq F_{d}(x,y^{*};u,v^{*})$.
+∎
 Theorem 3.4 (Envelope Inequality in Arbitrary Dimension).
 Let $A\in\mathbb{R}^{d_{1}\times d_{2}}_{>0}$ and $B\in\mathbb{R}^{d_{2}\times d_{3}}_{>0}$, where $d_{1},d_{2},d_{3}\geq 2$. Then
-$\boxed{\,R(AB)\;\leq\;\Phi\big(R(A),R(B)\big)\,}.$
+$R(AB)\leq\Phi\bigl(R(A),R(B)\bigr).$
 Proof.
-By definition, $R(AB)$ is the maximum distortion among all $2\times 2$ sub-blocks of $AB$. Choose rows $i,j$ of $A$ and columns $k,\ell$ of $B$ such that this maximum is attained.
-Let
+By definition, $R(AB)$ is attained by a $2\times 2$ sub-block of $AB$. Choose rows $i,j$ of $A$ and columns $k,\ell$ of $B$ attaining this maximum. Let
 […]
@@
 $\begin{pmatrix}x\cdot u&x\cdot v\\ y\cdot u&y\cdot v\end{pmatrix}$,
-whose distortion is
+and therefore
 $R(AB)=F_{d}(x,y;u,v).$
@@
 and
-$F_{d}(x,y;u,v)\;\leq\;F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime}).$
-Let $A^{\prime}$ be the $2\times 2$ matrix with rows $x^{\prime}$ and $y^{\prime}$, and let $B^{\prime}$ be the $2\times 2$ matrix with columns $u^{\prime}$ and $v^{\prime}$.
-Then $R(A^{\prime})=\operatorname{Dist}(x,y)=\alpha$ and $R(B^{\prime})=\operatorname{Dist}(u,v)=\beta$.
-Then
-$A^{\prime}B^{\prime}=\begin{pmatrix}x^{\prime}\!\cdot u^{\prime}&x^{\prime}\!\cdot v^{\prime}\\[2.0pt] y^{\prime}\!\cdot u^{\prime}&y^{\prime}\!\cdot v^{\prime}\end{pmatrix}$,
-and therefore
-$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})\;=\;R(A^{\prime}B^{\prime}).$
-Combining the above steps and using that $\Phi$ is increasing in each variable gives
+$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})\geq F_{d}(x,y;u,v).$
+Let $A^{\prime}$ be the $2\times 2$ matrix with rows $x^{\prime}$ and $y^{\prime}$, and let $B^{\prime}$ be the $2\times 2$ matrix with columns $u^{\prime}$ and $v^{\prime}$. Then
+$R(A^{\prime})=\operatorname{Dist}(x^{\prime},y^{\prime})=\alpha,\qquad R(B^{\prime})=\operatorname{Dist}(u^{\prime},v^{\prime})=\beta,$
+and
+$A^{\prime}B^{\prime}=\begin{pmatrix}x^{\prime}\!\cdot u^{\prime}&x^{\prime}\!\cdot v^{\prime}\\[2.0pt] y^{\prime}\!\cdot u^{\prime}&y^{\prime}\!\cdot v^{\prime}\end{pmatrix},\qquad\text{so}\qquad F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})=R(A^{\prime}B^{\prime}).$
+Using Theorem 2.2 and the monotonicity of $\Phi$,
 $R(AB)$ $=F_{d}(x,y;u,v)$
@@
 $=R(A^{\prime}B^{\prime})$
-$\leq\Phi(\alpha,\beta)$
-$\leq\Phi(R(A),R(B)).$
-This completes the proof.
+$\leq\Phi\bigl(R(A^{\prime}),R(B^{\prime})\bigr)$
+$=\Phi(\alpha,\beta)$
+$\leq\Phi\bigl(R(A),R(B)\bigr).$
 4 Distortion of Products
-Let $(A_{n})_{n\geq 1}$ be a sequence of positive matrices and define the partial products
-$P_{n}:=A_{n}\cdots A_{1}.$
-The matrices may have arbitrary dimensions, provided the products are well defined.
+Let $(A_{n})_{n\geq 1}$ be a sequence of positive matrices and set
+$P_{n}:=A_{n}\cdots A_{1},$
+assuming the products are well defined.
 It is convenient to work with the square root of the distortion,
@@
 $S(AB)\;\leq\;\Psi\bigl(S(A),S(B)\bigr),\qquad\Psi(p,q):=\frac{1+pq}{p+q}.$
-Since $\Psi$ is increasing in each variable, distortion bounds propagate along the product. If
+Since $\Psi$ is increasing, distortion bounds propagate along products. If
 $R(A_{n})\leq\alpha_{n},\qquad\text{i.e.}\qquad S(A_{n})\leq p_{n}:=\sqrt{\alpha_{n}},$
@@
 $q_{n+1}=\Psi(p,q_{n}),\qquad q_{1}=p.$
-With $p$ fixed, this iterates a single Möbius map. Writing
+This recursion is the iteration of a single Möbius map. Writing
 $\kappa:=\frac{p-1}{p+1},$
-its $n$th iterate can be expressed as
+we obtain the closed-form expression
 $q_{n}=1+\frac{2\,\kappa^{n}}{1-\kappa^{n}}.$
@@
 $R(P_{n})\leq q_{n}^{2}=1+\frac{4\,\kappa^{n}}{(1-\kappa^{n})^{2}}.$
-For convenience, we record the following standalone theorem.
+Collecting the preceding arguments, we record the following theorem.
 Theorem 4.1 (Finite-Product Distortion Bound).
@@
 In classical Birkhoff–Bushell (BB) theory one studies distortion on a log scale via the Hilbert metric
-$d_{H}(x,y)=\log\mathrm{Dist}(x,y),\qquad x,y\in\mathbb{R}^{n}_{>0}.$
+$d_{H}(x,y):=\log\mathrm{Dist}(x,y),\qquad x,y\in\mathbb{R}^{n}_{>0}.$
 If $B$ is an $n\times 2$ matrix with columns $x$ and $y$, and $A$ is any positive $n\times n$ matrix, then
 $d_{H}(x,y)=\log R(B),\qquad d_{H}(Ax,Ay)=\log R(AB).$
-In this log setting, our envelope inequality becomes
+In this logarithmic setting, the envelope inequality takes the form
 $d_{H}(Ax,Ay)\;\leq\;\Theta\!\left(d_{H}(x,y)\right),$
 where
-$\Theta(h)=2\log\!\left(\frac{1+p\,e^{h/2}}{\,p+e^{h/2}}\right),\qquad p=\sqrt{R(A)}.$
-In the nontrivial case $p>1$, a direct computation shows that $\Theta$ is strictly increasing, strictly concave, and satisfies
+$\Theta(h):=2\log\!\left(\frac{1+p\,e^{h/2}}{\,p+e^{h/2}}\right),\qquad p:=\sqrt{R(A)}.$
+A direct computation gives
+$\Theta^{\prime}(h)=\frac{(p^{2}-1)e^{h/2}}{(1+pe^{h/2})(p+e^{h/2})},\qquad\Theta^{\prime\prime}(h)=-\frac{(p^{2}-1)e^{h/2}p(e^{h}-1)}{2(1+pe^{h/2})^{2}(p+e^{h/2})^{2}}.$
+In the nontrivial case $p>1$, it follows that $\Theta$ is increasing and concave on $0,\infty)$, and satisfies
 $\Theta(0)=0,\qquad\Theta^{\prime}(0)=\frac{p-1}{p+1}=:\kappa(A).$
-Thus $\Theta$ lies below its tangent line at the origin:
+By concavity, $\Theta$ lies below its tangent line at the origin,
 $\Theta(h)\;\leq\;\kappa(A)\,h,$
-which yields the classical BB contraction:
+which yields the classical BB contraction estimate
 $d_{H}(Ax,Ay)\;\leq\;\kappa(A)\,d_{H}(x,y).$
-Note that
+Moreover,
 $\lim_{h\to\infty}\Theta(h)=\log R(A),$
-so the envelope bound $\Theta(h)$ remains uniformly bounded, whereas the BB linear bound
-$\kappa(A)\,h$
-diverges.
-Thus both bounds agree to first order at $h=0$ (and therefore yield the same Lyapunov–gap asymptotics for distortion of high powers of $A$), but for large $h$ the BB estimate increasingly overestimates the worst–case contraction allowed by $R(A)$, while the envelope bound remains sharp in the non-asymptotic regime.
-For illustration, Figure 2 compares the nonlinear envelope map $h\mapsto\Theta(h)$ with the classical BB line $k(A)h$, together with the saturation level $\log R(A)$.
+so the envelope bound remains uniformly bounded, whereas the linear BB bound $\kappa(A)h$ diverges.
+The two bounds therefore agree to first order at $h=0$, but for large $h$ the BB estimate increasingly overestimates the worst–case contraction allowed by $R(A)$, while the envelope bound remains sharp for all $h$.
+Figure 2 compares the nonlinear envelope $h\mapsto\Theta(h)$ with the classical BB line $\kappa(A)h$ and the saturation level $\log R(A)$.
 […]
 Figure 2: Comparison of the envelope $\Theta(h)$ with the classical Birkhoff–Bushell contraction bound $\kappa(A)h$ for $R(A)=9$.
-The envelope contraction saturates at $\log R(A)$, while the BB line grows without bound, showing that the BB estimate increasingly overestimates the worst–case contraction allowed by $R(A)$ for large $h$.
-Historical perspective.
-Birkhoff’s fundamental insight was that positive operators act as contractions in the Hilbert metric, allowing the Frobenius–Perron alignment phenomenon to be deduced from the Banach fixed-point theorem.
-Bushell later refined this viewpoint by computing an explicit contraction constant.
-In most contraction theorems such constants are technically useful but algebraically unremarkable: they serve as quantitative bounds and carry little intrinsic meaning.
-Bushell’s constant is a striking exception.
-Its closed form,
-$\kappa(A)=\frac{\sqrt{R(A)}-1}{\sqrt{R(A)}+1},$
-is unexpectedly symmetric and rigid—far too elegant, in retrospect, to be merely a sharp Lipschitz coefficient.
-Moreover, the metric framework of the Hilbert metric plays no essential role in the classical proof: it is used only to invoke the Banach fixed-point theorem, not to explain the mechanism of alignment itself.
-This indicates that the contraction phenomenon behind the Frobenius–Perron theory is not inherently metric but fundamentally algebraic, encoded in the behaviour of ratios rather than in the geometry induced by the Hilbert metric.
-This motivates our approach: the envelope inequality shows that worst-case distortion evolves under Möbius iterations with a common attracting fixed point at $1$.
-In this formulation, alignment arises from the explicit fixed-point dynamics of these maps rather than from a geometric contraction in the Hilbert metric.
-6 Discussion and outlook
-The envelope inequality is dimension-free, and the same collapse mechanism extends to a broad class of infinite matrices and positive operators on cones.
-The operator-theoretic generalizations will be developed elsewhere.
-A second direction concerns other projective quantities. Since the collapse argument in Section 3 uses only the four-point ratios
-$\frac{(x\!\cdot\!u)(y\!\cdot\!v)}{(x\!\cdot\!v)(y\!\cdot\!u)},$
-it is natural to conjecture that analogous two-dimensional collapse mechanisms may apply to a broader class of projective functionals.
-The envelope inequality also suggests applications in Markov chain theory.
-Classical asymptotic mixing bounds typically rely on Doeblin, Dobrushin, or Hilbert-metric methods, whereas the envelope provides an exact non-asymptotic contraction profile.
-This may offer an alternative route to quantitative bounds for finite-time mixing of positive stochastic matrices.
-The recursion in Section 4 provides a deterministic model for worst-case distortion in products of positive matrices; its transfer-matrix form makes the evolution completely explicit.
-It may also serve as a foundation for combining our methods with probabilistic techniques to study non-asymptotic behavior of random products.
-These directions lie beyond the scope of the present paper. Our purpose here is purely foundational: to isolate and prove the core envelope inequality in its simplest form.
-Appendix A Calculus Tools
-In this appendix we prove properties of the functions (for $\alpha,\beta\geq 1$)
-$f_{\alpha,\beta}(t)=\frac{(1+t)(1+\alpha\beta\,t)}{(1+\alpha t)(1+\beta t)},\qquad t>0,$
-and
-$\Phi(\alpha,\beta)=\left(\frac{1+\sqrt{\alpha\beta}}{\sqrt{\alpha}+\sqrt{\beta}}\right)^{2}.$
-Lemma A.1.
-The function $f_{\alpha,\beta}$ has the following properties:
-  1. 1.
-    $f_{\alpha,\beta}(t)\geq 1$ for all $t>0$;
-  2. 2.
-    if $\alpha,\beta>1$, the unique global maximum is attained at
-    $t^{*}=\frac{1}{\sqrt{\alpha\beta}},$
-    and
-    $f_{\alpha,\beta}(t)\leq f_{\alpha,\beta}(t^{*})=\Phi(\alpha,\beta);$
-  3. 3.
-    if $\alpha=1$ or $\beta=1$, then $f_{\alpha,\beta}(t)\equiv 1$.
-Proof.
-(1) Expanding $f_{\alpha,\beta}(t)\geq 1$ gives
-$(1+t)(1+\alpha\beta t)\;\geq\;(1+\alpha t)(1+\beta t),$
-which simplifies to
-$(\alpha-1)(\beta-1)t\geq 0,$
-valid for all $\alpha,\beta\geq 1$ and $t>0$.
-(2) Set $\varphi(t)=\log f_{\alpha,\beta}(t)$. A straightforward algebraic simplification yields
-$\varphi^{\prime}(t)=\,\frac{(\alpha-1)(\beta-1)\,(1-\alpha\beta t^{2})}{(1+t)(1+\alpha t)(1+\beta t)(1+\alpha\beta t)}.$
-Thus $\varphi^{\prime}(t)=0$ exactly at
-$t^{*}=\frac{1}{\sqrt{\alpha\beta}},$
-and this is the unique global maximum. Substituting $t^{*}$ gives the stated value of $f_{\alpha,\beta}(t^{*})$.
-(3) If $\alpha=1$ or $\beta=1$, the formula reduces to $f_{\alpha,\beta}(t)\equiv 1$. ∎
-Lemma A.2.
-For $\alpha,\beta\geq 1$, the envelope function $\Phi(\alpha,\beta)$ is increasing in each variable: for all $\alpha_{1}\leq\alpha_{2}$ and $\beta_{1}\leq\beta_{2}$,
-$\Phi(\alpha_{1},\beta)\leq\Phi(\alpha_{2},\beta),\qquad\Phi(\alpha,\beta_{1})\leq\Phi(\alpha,\beta_{2}).$
-Proof.
-Write $p=\sqrt{\alpha}$ and $q=\sqrt{\beta}$, so $p,q\geq 1$, and define
-$\psi(p,q)=\frac{1+pq}{p+q},$
-so that $\Phi(\alpha,\beta)=\psi(p,q)^{2}$. For fixed $q\geq 1$ we compute
-$\frac{\partial\psi}{\partial p}=\frac{q(p+q)-(1+pq)}{(p+q)^{2}}=\frac{q^{2}-1}{(p+q)^{2}}\;\geq\;0.$
-Thus $\psi$ is increasing in $p$, and therefore so is $\Phi=\psi^{2}$. By symmetry the same argument applies to $q$. ∎
+The envelope saturates at $\log R(A)$, while the BB bound grows without bound.